OpenStudy (anonymous):

What are the possible rational zeros of f(x) = x4 + 2x3 - 3x2 - 4x - 12? Can some one show me how to do this?

OpenStudy (alekos):

First you can try x= -1,0,1,2 I reckon that x = 2 will give you f(x) = 0 After that you can do a long division on the original quartic using x-2 as the divisor

OpenStudy (ccswims):

I would recommend using the Quadratic Formula for a big polynomial function like that one

OpenStudy (anonymous):

use long division

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-2)(x^3+4x^2+5x+6)=0

OpenStudy (anonymous):

do the same thing again

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-2)(x+3)(x^2+x+2)

OpenStudy (anonymous):

the last one use : \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

ax^2 +bx +c so a=1, b=1, c=2 \[\frac{ 1\pm \sqrt{1-4*1*2} }{ 2*1 }=\frac{ 1\pm \sqrt{1-8} }{ 2 }=\frac{ 1 }{ 2 }\pm \frac{ \sqrt{-7} }{2 }\]

OpenStudy (anonymous):

these roots are not real

OpenStudy (anonymous):

rational zeros

OpenStudy (alekos):

Ccs. Quadratic formula doesn't work for a quartic

OpenStudy (alekos):

Well done AJ. That's two real roots